### v2

# Jean DELHAYE

Abstract.
In this study, we construct an analog of the Brownian motion on free unitary quantum groups and compute its cutoff profile.
1. Introduction
Let $(X_{N})_{N\in\mathbb{N}}$ be a sequence of irreducible, aperiodic, finite-state Markov chains, with $\mu_{N}(t)$ denoting the distribution of $X_{N}$ after $t$ steps and $\mu_{N}(\infty)$ its stationary distribution. Define the total variation distance to equilibrium at time $t$ as
$d_{N}(t)=d_{\mathrm{TV}}(\mu_{N}(t),\mu_{N}(\infty)),$
where the total variation distance $d_{\mathrm{TV}}(\mu,\nu)$ between two probability measures $\mu$ and $\nu$ on a compact set $K$ is given by
$d_{\mathrm{TV}}(\mu,\nu)=\sup_{A\subset K}|\mu(A)-\nu(A)|,$
where the supremum is taken over all Borel sets. Consider a sequence of times $(t_{N})_{N\in\mathbb{N}}$. We say that $(X_{N})_{N\in\mathbb{N}}$ exhibits a cutoff in total variation distance at time $(t_{N})_{N\in\mathbb{N}}$ if, for every $\epsilon>0$,
$\lim_{N\to\infty}d_{N}((1-\epsilon)t_{N})=1\quad\text{and}\quad\lim_{N\to\infty}d_{N}((1+\epsilon)t_{N})=0.$
This indicates that the convergence to equilibrium undergoes a sharp transition, collapsing from $1$ to 00 around the time $t_{N}$.
The first example of the cutoff phenomenon in groups comes from the work of P. Diaconis and M. Shahshahani in the 1980s on finite groups [15]. This groundbreaking discovery sparked significant interest in the phenomenon, leading to the identification of numerous examples across various contexts. Notable cases include Brownian motion on simple compact Lie groups and symmetric spaces [28], as well as random walks on random graphs [26]. In this paper, we focus on quantum compact groups, an object initially introduced by S.L. Woronowicz in [36]. The representation theory of compact groups extends naturally to the quantum setting, enabling the application of similar techniques. This has been studied extensively, starting with finite quantum groups, as explored by J.P. McCarthy in [27], and extended to infinite compact quantum groups in works such as [20, 21].
To gain deeper insight into this phenomenon, we zoom in on the interval where this sharp transition occurs. The cutoff phenomenon suggests that the width of this interval is negligible compared to the sequence $(t_{N})_{N\in\mathbb{N}}$, prompting us to investigate the next significant higher-order term. Specifically, if there exists a sequence $(s_{N})_{N\in\mathbb{N}}$ and a continuous function $f$ that decreases from $1$ to 00 such that, for all $c\in\mathbb{R}$,
$\lim_{N\to\infty}d_{N}(t_{N}+cs_{N})=f(c),$
then $f$ is referred to as the cutoff profile or limit profile of $(X_{N})_{N\in\mathbb{N}}$.
While determining a cutoff profile is inherently challenging, there are several processes for which such profiles have been established. For instance, in the case of discrete processes like random walks on finite groups, the random transposition shuffle [33] exhibits the following asymptotic behavior:
$d_{N}\left(\frac{N}{2}\ln N+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Pois}(1+e^{-c}),\mathrm{Pois}(1)\right),\quad c\in\mathbb{R}.$
Other examples include the lazy random walk on the hypercube [31], the dovetail shuffle [5], the simple exclusion process on the circle [24], the Ehrenfest Urn model with multiple urns [30], the Gibbs Sampler [30], and random cycles [30].
The study of cutoff profiles on compact quantum groups is relatively recent and primarily based on the work [23], where the authors compute the profile for the Brownian motion on the quantum orthogonal group $O_{N}^{+}$ and the quantum permutation group $S_{N}^{+}$, as well as for quantum random transpositions on $S_{N}^{+}$.
While the Brownian motion is the most natural continuous process on simple compact Lie groups, defining a clear analog on most compact quantum groups is not straightforward. The orthogonal and permutation quantum groups provide natural processes that can be interpreted as such (stemming from analogies between the central generating functionals on $O_{N}^{+}$ and $S_{N}^{+}$, as proved in [12, Thm 10.2] and [18, Thm 10.10], and the classical case for compact Lie groups in [25]). However, extending this concept to other compact quantum groups presents significant challenges.
Our main result is the identification and analysis of the Brownian motion on the unitary quantum group $U_{N}^{+}$, which is detailed in Section 3, along with the partial computation of its limit profile, presented in Section 5. Specifically, we establish in Corollary 5.2 that, for appropriate extensions $\widetilde{d}_{N}$ of the distances $d_{N}$ in the quantum setting, we have
$\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),\quad c\geq 0,$
$\underset{N\to\infty}{\lim\sup}\,\widetilde{d}_{N}\left(N\ln(\sqrt{2}N)+cN\right)\geq d_{\mathrm{TV}}\left(\mathrm{Meix}^{+}(-e^{-c},0)*\delta_{e^{-c}},\nu_{\mathrm{SC}}\right),\quad c<0,$
where $\nu_{\mathrm{SC}}$ denotes the semicircle distribution and $\mathrm{Meix}^{+}$ refers to the free Meixner distribution (see Subsection 2.2 for definitions).
In most cases, the cutoff profile typically takes the form
$c\longmapsto d_{\mathrm{TV}}\left(\nu_{c},\mathrm{Haar}_{\infty}\right),\quad c\in\mathbb{R},$
where $\nu_{c}$ represents a limit measure (or state) for the process at time $t_{N}+cs_{N}$. The main difficulty in the quantum setting arises because the process is not absolutely continuous with respect to the stationary distribution whenever $c<c_{0}$ (for some $c_{0}\in(0,1)$, depending on the scaling of the sequences). This phenomenon, first noted in [23], shows that processes exhibit an atom in certain regions. In our case, the central algebra of the quantum unitary group $U_{N}^{+}$ is not commutative, unlike $O_{N}^{+}$ and $S_{N}^{+}$. Additionally, in the region where absolute continuity is lost, the singular part does not reduce to a single atom, but rather involves a more complex structure that we were unable to fully identify, which is why we can only establish an lower bound in this region.
The article concludes in Section 6, where we discuss the general observation that cutoff profiles tend to be similar across different structures. This raises the question of whether the structure alone determines the cutoff profile. We address this by providing an example of a process on $O_{N}^{+}$ that exhibits a distinct limit profile from that of Brownian motion.
Acknowledgement
We express our deepest gratitude to our advisor, A. Freslon, for proposing this problem and for his invaluable guidance throughout this work. We are also sincerely thankful to A. Skalski and L. Teyssier for their insightful feedback on earlier versions of this article, which helped us address inconsistencies and clarify certain points.
For more information, we encourage the reader to consult [13] for an excellent source on mixing times, and [22], [29], or [4] for a detailed study of compact quantum groups and their representation theory.
2. Preliminaries
In this paper, we will focus on compact quantum groups, which may be unfamiliar to probabilist readers. To address this, we will dedicate a preliminary section to defining and presenting key aspects. Specifically, we will introduce free unitary and orthogonal quantum groups, along with relevant results regarding Lévy processes on these objects.
2.1. Free unitary quantum groups
Free unitary quantum groups, a specific class of compact quantum groups, were introduced by S. Wang in [35]. Quantum groups, more broadly, were originally defined by S.L. Woronowicz [36]. These structures involve C*-algebras, fitting their noncommutative topological nature. However, in this article, we will present an alternative definition that emphasizes the algebraic aspects, making it more accessible to non-expert readers.
Let us recall that a *-algebra is a complex unital algebra $A$ endowed with an involution $A\to A,x\mapsto x^{*}$, i.e. an antimultiplicative antilinear map such that $x^{**}=x$ for all $x\in A$. Also, a *-ideal $B$ of $A$ is a *-subalgebra such that $aB\cup Ba\subset B$ for all $a\in A$.
Definition 2.1.
We define $\mathcal{O}(U_{N}^{+})$ to be the universal *-algebra generated by $N^{2}$ elements $\{u_{ij}\}_{1\leq i,j\leq N}$ such that
$\sum_{k=1}^{N}u_{ik}u_{jk}^{*}=\delta_{ij}=\sum_{k=1}^{N}u_{ki}^{*}u_{kj}\sum_{k=1}^{N}u_{ik}^{*}u_{jk}=\delta_{ij}=\sum_{k=1}^{N}u_{ki}u_{kj}^{*}.$
In other words, $\mathcal{O}(U_{N}^{+})=\mathbb{C}\langle u_{ij}\rangle_{1\leq i,j\leq N}/I$ where $\mathbb{C}\langle u_{ij}\rangle_{1\leq i,j\leq N}$ denotes the *-algebra of noncommutative polynomials in variables $u_{ij}$, $u_{ij}^{*}$ ($1\leq i,j\leq N$) and $I$ denotes the *-ideal generated by the elements
$\left\{\sum_{k=1}^{N}u_{ik}u_{jk}^{*}-\delta_{ij},\sum_{k=1}^{N}u_{ki}^{*}u_{kj}-\delta_{ij},\sum_{k=1}^{N}u_{ik}^{*}u_{jk}-\delta_{ij},\sum_{k=1}^{N}u_{ki}u_{kj}^{*}-\delta_{ij}\right\}.$
Let $U_{N}$ be the usual unitary group, $c_{ij}:U_{N}\to\mathbb{C}$ the function sending a matrix to its $(i,j)$-th coefficient and let $\mathcal{O}(U_{N})$ be the algebra of regular functions on $U_{N}$, i.e. the *-algebra generated by the functions $c_{ij}$, where the involution corresponds to the complex conjugation: $c_{ij}^{*}=\overline{c}_{ij}$. Then, quotienting $\mathcal{O}(U_{N}^{+})$ by its commutator ideal yields a surjection
$\pi\left\{\begin{matrix}\mathcal{O}(U_{N}^{+})&\longrightarrow&\mathcal{O}(U_{N})\\ u_{ij}&\longmapsto&c_{ij}.\end{matrix}\right.$
In that sense, $U_{N}^{+}$ can be thought of as a noncommutative version of $U_{N}$. The group structure of $U_{N}$ can be recovered in an algebraic way on $\mathcal{O}(U_{N})$ via the following remark:
$c_{ij}(gh)=\sum_{k=1}^{N}c_{ik}(g)c_{kj}(h)=\sum_{k=1}^{N}(c_{ik}\otimes c_{kj})(g,h),\quad g,h\in U_{N},1\leq i,j\leq N.$
The “group law” on $U_{N}^{+}$ is therefore given by the unique *-homomorphism $\Delta:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})\otimes\mathcal{O}(U_{N}^{+})$, called coproduct, such that
$\Delta(u_{ij})=\sum_{k=1}^{N}u_{ik}\otimes u_{kj}.$
The existence of $\Delta$ follows from the universal property. Additionally, let us outline the existence of the counit and the antipode, which are crucial objects. The counit is given by the homomorphism $\varepsilon:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$, defined by $\varepsilon(u_{ij})=\delta_{ij}$. The antipode, on the other hand, is the antihomomorphism $S:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})$, defined by $S(u_{ij})=u_{ji}^{*}$. These maps serve as the unit and the inverse map in the noncommutative setting, respectively.
In this context, probability measures can be generalized by identifying them with their integration linear forms. These forms correspond to states, which are unital positive linear forms on $\mathcal{O}(U_{N}^{+})$. Notably, there exists a particular state that serves as the analog of the Haar measure on $U_{N}^{+}$ (see [36]).
Theorem 2.2.
There is a unique state $h$ on $U_{N}^{+}$ such that
$(\operatorname{id}\otimes h)\circ\Delta(x)=h(x)\otimes 1=(h\otimes\operatorname{id})\circ\Delta(x),\quad x\in\mathcal{O}(U_{N}^{+}).$
It is called the Haar state of $U_{N}^{+}$.
Remark 2.3.
Let us point out the fact that the Haar state is tracial, i.e. we have $h(xy)=h(yx)$ for any $x,y\in\mathcal{O}(U_{N}^{+})$.
The use of representation theory as a powerful tool to investigate the asymptotic behavior of random walks on groups has been established since the seminal works of P. Diaconis and his coauthors (see, for example, [14, Chap 4]). To describe the representations of $U_{N}^{+}$, we will make use of the free orthogonal quantum group $O_{N}^{+}$, given by the *-algebra that is the quotient $\mathcal{O}(O_{N}^{+})\simeq\mathcal{O}(U_{N}^{+})/\{u_{ij}-u_{ij}^{*}\}$. We will denote by $o_{ij}$ the image of $u_{ij}$ under the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$. The coproduct factors through the quotient map yielding a compact quantum group structure.
If $z$ denotes the identity function on $\mathbb{T}$, then it was proven in [3, Prop 7] that the map $u_{ij}\mapsto o_{ij}z$ extends to an isomorphism of compact quantum groups between $\mathcal{O}(U_{N}^{+})$ and its image in $\mathcal{O}(O_{N}^{+})*\mathcal{O}(\mathbb{T})$. T. Banica proved in [2] that the characters of the irreducible representations of $O_{N}^{+}$ may be labelled by the nonnegative integers $(\chi_{n})_{n\in\mathbb{N}}$ such that
$\chi_{0}=1,\quad\chi_{1}=\sum_{j=1}^{N}o_{jj}\quad\hbox{\&}\quad\chi_{1}\chi_{n}=\chi_{n+1}+\chi_{n-1},\quad n\in\mathbb{N}^{*}.$
This relation is reminiscent of that of Chebyshev polynomials of the second kind. That is the sequence of polynomials $(P_{n})_{n\in\mathbb{N}}$ recursively given by
$P_{0}=1,\quad P_{1}=X\quad\hbox{\&}\quad XP_{n}=P_{n+1}+P_{n-1},\quad n\in\mathbb{N}^{*}.$
Note that we have the relation $\chi_{n}=P_{n}(\chi_{1})$ for all $n\in\mathbb{N}$.
It then follows from the description of the representation theory of free products [35] that the characters of $U_{N}^{+}$ can be recovered as products of characters of $O_{N}^{+}$ and powers of $z$. A precise description is given in [34, Prop 4.3] which we reproduce here.
Theorem 2.4.
The characters of the nontrivial irreducible representations of $U_{N}^{+}$ are the elements of the form
$\chi^{\epsilon}_{\mathbf{n}}:=z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{\epsilon_{2}}...z^{\epsilon_{p}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}},$
where $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, $\epsilon=\epsilon_{1}\in\{\pm 1\}$, $[\epsilon]_{-}=\min(\epsilon,0)$, $[\epsilon]_{+}=\max(\epsilon,0)$ and $\epsilon_{i+1}=(-1)^{n_{i}+1}\epsilon_{i}$. The dimensions can be recovered by applying the counit
$d_{\mathbf{n}}:=\varepsilon\big{(}z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{\epsilon_{2}}...z^{\epsilon_{p}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}}\big{)}=P_{n_{1}}(N)...P_{n_{p}}(N).$
We denote by $\mathcal{O}(U_{N}^{+})_{0}$ the algebra generated by the characters and call it the central algebra of $U_{N}^{+}$. An important feature of this subalgebra is the existence of a conditional expectation (see [1, Chap 9]) $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$ that leaves the Haar state invariant, let us recall the definition of such an object.
Definition 2.5.
Let $A$ be a $C^{*}$-algebra (that is a Banach *-algebra such that $\|a^{*}a\|=\|a\|^{2}$ for all $a\in A$) and $B\subset A$ a subalgebra of $A$. We call conditional expectation from $A$ onto $B$ a linear map $\mathbb{E}:A\to B$ satisfying:
  1. (i)
    $\mathbb{E}[A_{+}]\subset B_{+}$ where $A_{+}=\{a^{*}a:a\in A\}$ and $B_{+}=A_{+}\cap B$;
  2. (ii)
    $\mathbb{E}[b]=b$ for all $b\in B$;
  3. (iii)
    $\mathbb{E}[bab^{\prime}]=b\mathbb{E}[a]b^{\prime}$ for all $a\in A$ and $b,b^{\prime}\in B$.
Note that a conditional expectation is always a norm-one projection. The importance of such an object will become apparent later on.
Let us now describe the quantum analogue of a Lévy process on $U_{N}^{+}$. We call (quantum) Lévy process on $U_{N}^{+}$ any right-continuous convolution semigroup of states, i.e. a family $(\psi_{t})_{t\geq 0}$ of states on $U_{N}^{+}$ such that
  • $\psi_{0}=\varepsilon$;
  • $\psi_{t}\star\psi_{s}:=(\psi_{t}\otimes\psi_{s})\circ\Delta=\psi_{t+s}$;
  • $\psi_{t}\to\psi_{0}$ weakly as $t\to 0$.
We call generating functional on $U_{N}^{+}$ an hermitian functional $L:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ that vanishes on $1$ and is positive on $\ker\varepsilon$. There is a one-to-one correspondence between Lévy processes and generating functionals (see [19, Sec 1.5]) via the formulas
$L$ $=\lim_{t\to 0}\frac{\psi_{t}-\varepsilon}{t},$
$\psi_{t}$ $=\exp_{\star}(tL)=\sum_{n\geq 0}\frac{(tL)^{\star n}}{n!},\quad t\geq 0.$
Definition 2.6.
We will say that a Lévy process $(\psi_{t})_{t\geq 0}$ is central if all the $\psi_{t}$’s are central, i.e. if $\psi_{t}=\psi_{t}\circ\mathbb{E}$ for all $t\geq 0$. Equivalently, a Lévy process is central if its associated generating functional $L$ is central, that is, $L=L\circ\mathbb{E}$.
The notion of centrality will become relevant in Section 3 as we introduce the quantum unitary Brownian motion as a central Lévy process. Observe that for any generating functional $L$ on $\mathcal{O}(U_{N}^{+})$, the composition $L\circ\mathbb{E}$ also defines a generating functional on $\mathcal{O}(U_{N}^{+})$ which we call the centralized generating functional. Thus, we may consider centralized generating functional and when doing so have in mind that all their information is contained within the restriction of $L$ to the central algebra $\mathcal{O}(U_{N}^{+})_{0}$.
2.2. The cutoff phenomenon
In this subsection, we introduce the concept of the cutoff phenomenon within the framework of quantum groups. The cutoff phenomenon, a sharp transition in the convergence to equilibrium, is a significant topic in probability theory and has been widely studied in classical settings. Here, we aim to extend this understanding to the realm of quantum groups, particularly focusing on the unitary quantum group. We will provide necessary definitions, discuss the associated Lévy processes, and outline the framework used to investigate the limit profiles.
One may consider the universal enveloping C*-algebra $C(U_{N}^{+})$ associated with $\mathcal{O}(U_{N}^{+})$ (see, for instance, [7, Sec II.8.3]). By definition, any state on $\mathcal{O}(U_{N}^{+})$ uniquely extends to a state on $C(U_{N}^{+})$, thus yielding an element of the Fourier-Stieltjes algebra, which corresponds to the topological dual $C(U_{N}^{+})^{*}$ of $C(U_{N}^{+})$. By the Riesz representation theorem, this dual space can be viewed as a noncommutative analogue of the measure algebra, equipped with the total variation norm. We denote the norm on this dual space as $\|\cdot\|_{FS}$, referring to it as the Fourier-Stieltjes norm.
Furthermore, the topological double dual $C(U_{N}^{+})^{**}$ of $C(U_{N}^{+})$ is recognized as the universal enveloping von Neumann algebra of $C(U_{N}^{+})$, serving as a noncommutative and universal analogue of classical measure spaces. Using the theory of Haagerup’s noncommutative $\mathrm{L}^{p}$-spaces, we can readily adapt the argument from [21, Lem 2.6] to show that:
$\frac{1}{2}\|\varphi-\psi\|_{FS}=\sup_{p\in\mathcal{P}}|\varphi(p)-\psi(p)|,$
where $\mathcal{P}$ denotes the set of orthogonal projections in $C(U_{N}^{+})^{**}$, which can be thought of as the noncommutative counterparts to indicator functions of Borel subsets. The proof details are omitted here, as they are not required for our current discussion. This generalized total variation distance will be employed in our analysis of cutoff phenomena, and we will denote this distance as
$d_{\mathrm{TV}}(\varphi,\psi)=\frac{1}{2}\|\varphi-\psi\|_{FS}.$
In the classical setting, a particularly significant case arises when both $\mu$ and $\nu$ are absolutely continuous with respect to the Haar measure. We can consider a similar scenario in the quantum context. Define an inner product on $\mathcal{O}(U_{N}^{+})$ by $\langle x,y\rangle=h(xy^{*})$. Completing this inner product space yields the Hilbert space $\mathrm{L}^{2}(U_{N}^{+})$, and $\mathcal{O}(U_{N}^{+})$ embeds into $\mathrm{B}(\mathrm{L}^{2}(U_{N}^{+}))$ via left multiplication (see [29, Cor 1.7.5] and the comments thereafter). The weak closure of this image is denoted by $\mathrm{L}^{\infty}(U_{N}^{+})$ and forms a von Neumann algebra.
If $\varphi:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ is a linear map that extends to a normal bounded map on $\mathrm{L}^{\infty}(U_{N}^{+})$, then $\varphi$ is an element of the Fourier algebra, which is the Banach space predual $\mathrm{L}^{\infty}(U_{N}^{+})_{*}$ of $\mathrm{L}^{\infty}(U_{N}^{+})$. In this case, $\|\varphi\|_{FS}=\|\varphi\|_{\mathrm{L}^{\infty}(U_{N}^{+})_{*}}$ (see [11, Prop 3.14]), which further implies, by [21, Lem 2.6], that:
$d_{\mathrm{TV}}(\varphi,\psi)=\sup_{p\in\widetilde{\mathcal{P}}}|\varphi(p)-\psi(p)|,$
where $\widetilde{\mathcal{P}}$ is the set of orthogonal projections in $\mathrm{L}^{\infty}(U_{N}^{+})$. It is important to note that for this formula to be applicable, the states $\varphi$ and $\psi$ must extend to the von Neumann algebra $\mathrm{L}^{\infty}(U_{N}^{+})$.
Definition 2.7.
Let $(\mathbb{G}_{N},\varphi^{(N)}_{t\geq 0})_{N\in\mathbb{N}}$ be a sequence of compact quantum groups each equipped with a Lévy process. We say that they exhibit a cutoff phenomenon at time $(t_{N})_{N\in\mathbb{N}}$ if for any $\epsilon>0$ we have
$d_{\mathrm{TV}}(\varphi_{t_{N}(1-\epsilon)}^{(N)},h_{N})\underset{N\to\infty}{\longrightarrow}1\quad\hbox{\&}\quad d_{\mathrm{TV}}(\varphi_{t_{N}(1+\epsilon)}^{(N)},h_{N})\underset{N\to\infty}{\longrightarrow}0.$
where $h_{N}$ denotes the Haar state of $\mathbb{G}_{N}$. More precisely, given a continuous function $f$ decreasing from $1$ to 00 such that
$d_{\mathrm{TV}}(\varphi_{t_{N}+cs_{N}},h)\underset{N\to\infty}{\longrightarrow}f(c),\quad c\in\mathbb{R},$
for some sequence $(s_{N})_{N\in\mathbb{N}}$ with $s_{N}=o(t_{N})$, we say that $f$ is the limit profile of the process.
Remark 2.8.
Note that a limit profile is unique only up to affine transformation. Specifically, suppose a given limit profile $f$ is attained along the sequences $(t_{N},s_{N})_{N}$. Then, for any $a>0$ and $b\in\mathbb{R}$, the limit profile $c\mapsto f(ac+b)$ is achieved along the sequences $(t_{N}+bs_{N},as_{N})_{N}$. Now, consider a single process that exhibits two distinct limit profiles, $f$ and $f^{\prime}$, along the sequences $(t_{N},s_{N})_{N}$ and $(t^{\prime}_{N},s^{\prime}_{N})_{N}$, respectively. Up to an affine transformation, we may assume that these limit profiles coincide at two points, say 00 and $1$, and that $f$ is strictly decreasing in a neighborhood of both 00 and $1$ (as the function decreases from 1 to 0). Then, if $d_{N}(t)$ denotes the distance between the process and the Haar state at time $t$, we have:
$\lim_{N\to\infty}d_{N}(t_{N})=f(0)=\lim_{N\to\infty}d_{N}(t^{\prime}_{N}),\lim_{N\to\infty}d_{N}(t_{N}+s_{N})=f(1)=\lim_{N\to\infty}d_{N}(t^{\prime}_{N}+s^{\prime}_{N}).$
Given that the $d_{N}$’s are decreasing (see [23, Lem 2.6]), it follows that $t^{\prime}_{N}=t_{N}+o(s_{N})$ and $s^{\prime}_{N}=s_{N}+o(s_{N})$. To see this, suppose, for instance, that the first equality does not hold. This would imply that, possibly after passing to a subsequence, we have
$t^{\prime}_{N}\geq t_{N}+cs_{N}$
for some $c>0$ and for $N$ sufficiently large (the inequality could be reversed and $c<0$, but the argument remains the same). This inequality would further imply that
$d_{N}(t_{N})\leq d_{N}(t_{N}+cs_{N})\underset{N\to\infty}{\longrightarrow}f(c)<f(0),$
which is a contradiction. A similar argument shows that the second equality must hold as well. Hence, we conclude that $f=f^{\prime}$.
In this paper, we explore the limit profile of the Brownian motion on $U_{N}^{+}$ which is shared by the Brownian motion on $O_{N}^{+}$, with the primary challenge being the identification of suitable sequences $(t_{N})_{N\in\mathbb{N}}$ and $(s_{N})_{N\in\mathbb{N}}$. We denote by $\nu_{\mathrm{SC}}$ the semi-circular distribution, which is defined as follows:
$\mathrm{d}\nu_{\mathrm{SC}}=\frac{\sqrt{4-x^{2}}}{2\pi}\mathbf{1}_{|x|<2}\mathrm{d}x,$
and its $\mathrm{L}^{2}$-space has the Chebyshev polynomials as an orthonormal basis, as introduced just before Theorem 2.4. For any real number $c$, let $\eta_{c}$ represent the distribution defined by
$\mathrm{d}\eta_{c}=(1-e^{2c})\mathbf{1}_{c<0}\delta_{e^{c}+e^{-c}}+\frac{e^{c}\sqrt{4-x^{2}}\mathbf{1}_{|x|<2}}{2\pi(e^{c}+e^{-c}-x)}\mathrm{d}x.$
In essence, $\eta_{c}$ is the shifted free Meixner law $\mathrm{Meix}(-e^{-c},0)*\delta_{e^{-c}}$. For further details on free Meixner laws, we refer the reader to [8, Sec 2.2]. It is important to note that the semi-circular distribution can be considered a free Meixner law with parameters $(0,0)$, emphasizing that $\eta_{c}$ is effectively a deformation of $\nu_{\mathrm{SC}}$. Moreover, it has been established in [23, Lem 3.12] that $\eta_{c}$ is the unique measure satisfying
$\eta_{c}(P_{n})=e^{-cn},\quad n\in\mathbb{N}.$
The limit profile of the Brownian motion on $U_{N}^{+}$ is determined by the function mapping each $c\in\mathbb{R}$ to the total variation distance between $\eta_{c}$ and $\nu_{\mathrm{SC}}$.
3. The quantum unitary Brownian motion
3.1. Defining the quantum unitary Brownian motion
It is not entirely clear which generating functional on $U_{N}^{+}$ could serve as the analogue of the Brownian motion. Fortunately, a well-defined concept of Brownian motion on $O_{N}^{+}$ has been established. F. Cipriani, U. Franz, and A. Kula provided a classification of central generating functionals on $O_{N}^{+}$ in [12, Thm 10.2] using an analogue of the Lévy-Khinchine formula. Specifically, they determined that all such functionals have the following form:
(3.1) $\chi_{n}\longmapsto-bP^{\prime}_{n}(N)-\int_{-N}^{N}\frac{P_{n}(N)-P_{n}(x)}{N-x}\mathrm{d}\nu(x),$
for some $b\geq 0$ and positive Borel measure $\nu$ on $[-N,N)$.
Comparing this formula with the one established by M. Liao for classical compact Lie groups in [25], we see that the process corresponding to $\nu=0$ plays a role analogous to the one associated to the Laplace-Beltrami operator. As a consequence, this functional is called the Brownian motion on $O_{N}^{+}$.
We lack a similar decomposition for central generating functionals on $U_{N}^{+}$, making it challenging to define Brownian motion on this quantum group. Fortunately, the Brownian motion on $O_{N}^{+}$ has other properties that we can leverage to define a Brownian motion on $U_{N}^{+}$.
Firstly, any centralized Gaussian generating functional on $O_{N}^{+}$ qualifies as a Brownian motion (see the Appendix for a proof). Secondly, as proved in [10, Prop 3.9], the Laplace-Beltrami operator on $O_{N}$ acts as a Brownian motion on $O_{N}^{+}$ when composed with the quotient map $\mathcal{O}(O_{N}^{+})\to\mathcal{O}(O_{N})$.
[…]
$U_{w}:=U_{w_{1}}\otimes...\otimes U_{w_{n}}\quad\text{and}\quad\chi_{w}:=\chi_{U_{w}},$
where $U_{\lozenge}=U$ and $U_{\blacklozenge}=\overline{U}$. We also denote by $p(w)$, $q(w)$, and $\ell(w)$ the number of occurrences of $\lozenge$, $\blacklozenge$, and the total number of letters in the word $w$, respectively. We claim that functionals $L:\mathcal{O}(U_{N}^{+})_{0}\to\mathbb{C}$, satisfying for any word $w$:
(3.2) $L_{w}:=L(\chi_{w})=-\ell N^{\ell-1}\alpha-\Big{(}{(p-q)^{2}-\ell}\Big{)}N^{\ell-2}\beta,$
where $p:=p(w)$, $q:=q(w)$, and $\ell:=\ell(w)$, for some $\alpha,\beta\geq 0$ with $\alpha\geq\beta/N$, adhere to the expected behavior and properties of Brownian motions, thereby justifying their classification as such.
As expected, centralized real-valued Gaussian generating functionals on $U_{N}^{+}$ are of this form (see the Appendix for a proof). Let us prove that the Laplace-Beltrami operator on $U_{N}$ is of this form when considered on $U_{N}^{+}$.
The Lie algebra $\mathfrak{u}_{N}$ of the unitary group $U_{N}$ consists of $N\times N$ skew-Hermitian matrices. The family
$X_{ij}=\left\{\begin{array}[]{cc}{E_{ij}-E_{ji}/2\sqrt{N}}&\text{if }i<j,\\ {\mathrm{i}E_{ii}/\sqrt{2(N-1)}}&\text{if }i=j,\\ {\mathrm{i}(E_{ij}+E_{ji})/2\sqrt{N}}&\text{if }i>j,\end{array}\right.\quad 1\leq i,j\leq N,$
where $E_{ij}$ is the matrix with a 1 in the $(i,j)$-th entry and zeros elsewhere, forms an orthonormal basis for the negative Killing form of $\mathfrak{u}_{N}$, defined by
$B(X,Y)=-2N\operatorname{Tr}(XY)+2\operatorname{Tr}(X)\operatorname{Tr}(Y).$
We will use $X_{ij}$ to refer to both the matrix and the associated differential operator. The Laplace-Beltrami operator on $U_{N}$ is defined as $L=(\sum_{i,j}X_{ij}^{2})\big{|}_{g=e}$. We will always consider $L$ through the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N})$ and conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})_{0}\to\mathcal{O}(U_{N}^{+})$.
Proposition 3.1.
[…]
$L(\chi_{w})=-\ell(w)N^{\ell(w)-1}\alpha-\Big{(}(p(w)-q(w))^{2}-\ell(w)\Big{)}N^{\ell(\beta)-2}\beta,$
with $\alpha=(N^{2}-N+1)/2(N-1)$ and $\beta=N/2(N-1)$.
Proof.
Let us compute.
$L_{\circ}:=L(\chi_{U})$ $=\sum_{i,j}X_{ij}^{2}\operatorname{Tr}|_{e}=\sum_{i,j}\frac{\partial^{2}}{\partial s\partial t}\left.\operatorname{Tr}\left(\exp(sX_{ij})\exp(tX_{ij})\right)\right|_{t,s=0}$
$=\sum_{i,j}\operatorname{Tr}(X_{ij}^{2})=-\frac{N^{2}-N+1}{2(N-1)}=-\alpha;$
$L_{\circ\circ}:=L(\chi_{U\otimes U})$ $=\sum_{i,j}X_{ij}^{2}(\operatorname{Tr}\cdot\operatorname{Tr})|_{e}$
$=\sum_{i,j}X_{ij}\left.\left(X_{ij}(\operatorname{Tr})\operatorname{Tr}+\operatorname{Tr}X_{ij}(\operatorname{Tr})\right)\right|_{e}$
$=\sum_{i,j}\left.\left(X_{ij}^{2}(\operatorname{Tr})\operatorname{Tr}+2X_{ij}(\operatorname{Tr})X_{ij}(\operatorname{Tr})+\operatorname{Tr}X_{ij}^{2}(\operatorname{Tr})\right)\right|_{e}$
$=2NL(\chi_{U})+2\sum_{i}\operatorname{Tr}(X_{ii})^{2}=2N\alpha-\frac{N}{N-1}$
$=-2N\alpha-2\beta;$
$L_{\circ\bullet}:=L(\chi_{U\otimes\overline{U}})$ $=\sum_{i,j}X_{ij}^{2}(\operatorname{Tr}\cdot\overline{\operatorname{Tr}})|_{e}=\sum_{i,j}X_{ij}\left.\left(X_{ij}(\operatorname{Tr})\overline{\operatorname{Tr}}+\operatorname{Tr}X_{ij}(\overline{\operatorname{Tr}})\right)\right|_{e}$
$=\sum_{i,j}\left.\left(X_{ij}^{2}(\operatorname{Tr})\overline{\operatorname{Tr}}+2X_{ij}(\operatorname{Tr})X_{ij}(\overline{\operatorname{Tr}})+\operatorname{Tr}X_{ij}^{2}(\overline{\operatorname{Tr}})\right)\right|_{e}$
$=N(L(\chi_{U})-L(\chi_{\overline{U}}))+2\sum_{i,j}\operatorname{Tr}(X_{ij})\overline{\operatorname{Tr}}(X_{ij})=-2N\alpha+\frac{N}{N-1}$
$=-2N\alpha+2\beta.$
where $\overline{\operatorname{Tr}}$ denotes the function
$\overline{\operatorname{Tr}}:U_{N}\longrightarrow\mathbb{C},g\longmapsto\overline{\operatorname{Tr}(g)}.$
This shows the result for words of length $\leq 2$. Assume now that $w=w_{1}...w_{\ell}$ is a word of length $\ell\geq 3$ and set $p:=p(w)$ and $q:=q(w)$. Let us fix some notations. We will write $W$ to denote the map $\operatorname{Tr}_{w_{1}}\cdot...\cdot\operatorname{Tr}_{w_{n}}$ where $\operatorname{Tr}_{\lozenge}=\operatorname{Tr}$ and $\operatorname{Tr}_{\blacklozenge}=\overline{\operatorname{Tr}}$. We will also write $W_{n}$ (resp. $W_{mn}$) to denote the product map of the $\operatorname{Tr}_{w_{j}}$’s omitting $\operatorname{Tr}_{w_{n}}$ (resp. omitting $\operatorname{Tr}_{w_{m}}$ and $\operatorname{Tr}_{w_{n}}$). Let us finally compute.
$L(\chi_{w})$ $=\sum_{i,j}X_{ij}^{2}W|_{e}=\sum_{i,j}X_{ij}\left.\left(\sum_{n}X_{ij}(\operatorname{Tr}_{w_{n}})W_{n}\right)\right|_{e}$
$=\sum_{i,j}\left.\left(\sum_{n}X_{ij}^{2}(\operatorname{Tr}_{w_{n}})W_{n}+\sum_{m\neq n}X_{ij}(\operatorname{Tr}_{w_{m}})X_{ij}(\operatorname{Tr}_{w_{n}})W_{mn}\right)\right|_{e}$
$=-\sum_{n}W_{n}(e)\psi_{w_{n}}+\sum_{i,j}\sum_{m\neq n}W_{mn}(e)\operatorname{Tr}_{w_{m}}(X_{ij})\operatorname{Tr}_{w_{n}}(X_{ij})$
$=N^{\ell-1}(pL_{\lozenge}+qL_{\blacklozenge})-N^{\ell-2}\sum_{m\neq n}\frac{N(L_{w_{m}}+L_{w_{n}})-L_{w_{m}w_{n}}}{2}$
$=N^{\ell-2}\left(\frac{p(p-1)}{2}L_{\lozenge\lozenge}+\frac{q(q-1)}{2}L_{\blacklozenge\blacklozenge}+pqL_{\lozenge\blacklozenge}\right)-N^{\ell-1}(\ell-2)\left(pL_{\lozenge}+qL_{\blacklozenge}\right)$
$=-{\ell}N^{\ell-1}\alpha-\Big{(}{(p-q)^{2}-\ell}\Big{)}N^{\ell-2}\beta.$
This finally proves the result. ∎
This further strengthens the justification for labeling functionals of the form (3.2) as Brownian motions.
We are not asserting that every pair $(\alpha,\beta)$ satisfying $\alpha\geq\beta/N\geq 0$ necessarily defines a central generating functional through Formula (3.2), although this is something we anticipate to be true. Instead, we are stating that all generating functionals of this form should be regarded as Brownian motions. The existence of such examples is guaranteed by the Brownian motion on $U_{N}$. Similarly, processes where $\alpha\geq 0$ and $\beta=0$ can be constructed by composing a Brownian motion on $O_{N}^{+}$ (a generating functional described by Formula (3.1) with $\nu=0$) with the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$ and the conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$. From these two types of examples, many more can be generated, as the set of all generating functionals forms a convex cone.
Definition 3.2.
Any central generating functional on $U_{N}^{+}$ of the form
$L_{\alpha,\beta}:\chi_{w}\longmapsto-{\ell(w)}N^{\ell(w)-1}\alpha-\Big{(}{(p(w)-q(w))^{2}-\ell(w)}\Big{)}N^{\ell(w)-2}\beta,\quad w\hbox{ word on }\ \{\lozenge,\blacklozenge\},$
for some $\alpha,\beta\geq 0$ is called Brownian motion of parameter $(\alpha,\beta)$. We will also call Brownian motion of parameter $(\alpha,\beta)$, the associated Lévy process.
3.2. Computing the values of the Brownian motion
Let $L=L_{\alpha,\beta}$ be a Brownian motion on $U_{N}^{+}$. The goal of this subsection is to compute the values of $L$ on all characters of irreducible representations. Note that $L$ coincides with a Gaussian generating functional on the central algebra. Therefore, we may apply Lemma A.3 to $L$ on elements of the central algebra. Our first goal is to understand its values on elementary elements and products of two elementary elements. Specifically, we aim to determine the quantities of the form
$\Phi_{m}^{\epsilon_{1}}:=L(z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}})\quad\hbox{\&}\quad\Phi_{mn}^{\epsilon_{1}\eta_{1}}:=L(z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}\chi^{n}_{1}z^{[\eta_{2}]_{+}})$
where $m,n\in\mathbb{N}$, $\epsilon_{1},\eta_{1}\in\{\pm 1\}$, $\epsilon_{2}=(-1)^{m+1}\epsilon_{1}$ and $\eta_{2}=(-1)^{n+1}\eta_{1}$.
Lemma 3.3.
Let $\epsilon_{1},\eta_{1}\in\{\pm 1\}$ and $m,n\in\mathbb{N}$, we have the formulas
(3.3) $\Phi_{m}^{\epsilon_{1}}=-{m}N^{m-1}\alpha-({\mathfrak{p}_{m}-m})N^{m-2}\beta,$
and
(3.4) $\Phi_{mn}^{\epsilon_{1}\eta_{1}}=-({m+n})N^{m+n-1}\alpha-\Big{(}{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}-(m+n)}\Big{)}N^{m+n-2}\beta$
where,
$\mathfrak{p}_{m}=\left\{\begin{array}[]{ll}0&\hbox{if }m\hbox{ is even}\\ 1&\hbox{else}\end{array}\right.\quad\hbox{\&}\quad\mathfrak{p}_{mn}=\left\{\begin{array}[]{ll}0&\hbox{if }m\hbox{ or }n\hbox{ is even}\\ 2&\hbox{else}\end{array}\right.$
Proof.
Let $w$ be the word such that $z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}}=\chi_{w}$. It is not hard to see that $w$ is a word of alternating letters of length $m$. Particularly, $\ell(w)=m$ and $d(w)=\mathfrak{p}_{m}=d(w)^{2}$. Now, it is clear that Formula (3.3) follows from Equation (3.2).
For the second formula, we will split the study according to the parity of $m$ and $n$. Let $w_{m}$ and $w_{n}$ be the words such that $z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}}=\chi_{w_{m}}$ and $z^{[\eta_{1}]_{-}}\chi^{n}_{1}z^{[\eta_{2}]_{+}}=\chi_{w_{n}}$ and set $w:=w_{m}w_{n}$ so that $\chi_{w}=z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}\chi^{n}_{1}z^{[\eta_{2}]_{+}}$.
  • If $m$ and $n$ are both even, then clearly $d(w_{m})=d(w_{n})=0$ and so $d(w)=0=\mathfrak{p}_{m}-\mathfrak{p}_{n}$, thus Formula (3.4) holds.
  • If $m$ is even and $n$ is odd, then $d(w_{m})=0$ and $d(w_{n})=1$ and so $d(w)=1=\mathfrak{p}_{n}-\mathfrak{p}_{m}$ which is what we wanted. Same holds if $m$ is odd and $n$ is even.
  • Now, if both $m$ and $n$ are odd, then $d(w_{m})=d(w_{n})=1$ and $d(w)\in\{0,2\}$. As previously stated, both $w_{m}$ and $w_{n}$ are words of alternating letters, so the most present letter in $w_{m}$ and $w_{n}$ are their first letter. This gives us the following equivalence relation
    $d(w)=2$ $\iff w_{m}\hbox{ and }w_{n}\hbox{ have the same first letter}$
    $\iff\epsilon_{1}=\eta_{1}.$
    This shows the last case.
We can now understand the Brownian motion on some irreducible characters.
Proposition 3.4.
Let $m,n\in\mathbb{N}$ and $\epsilon_{1},\eta_{1}\in\{\pm 1\}$, we have
(3.5) $\Psi_{m}^{\epsilon_{1}}:=L(z^{[\epsilon_{1}]_{-}}\chi_{m}z^{[\epsilon_{2}]_{+}})=-\left(\alpha-\frac{\beta}{N}\right)P_{m}^{\prime}(N)-\beta\frac{\mathfrak{p}_{m}}{N^{2}}P_{m}(N),$
and
(3.6) $\begin{split}\Psi_{mn}^{\epsilon_{1}\eta_{1}}:&=L(z^{[\epsilon_{1}]_{-}}\chi_{m}z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}\chi_{n}z^{[\eta_{2}]_{+}})\\ &=-\left(\alpha-\frac{\beta}{N}\right)(P_{m}P_{n})^{\prime}(N)-\beta\frac{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}}{N^{2}}(P_{m}P_{n})(N)\end{split}$
where $\epsilon_{2}=(-1)^{m+1}\epsilon_{1}$ and $\eta_{2}=(-1)^{n+1}\eta_{1}$.
Proof.
Let us write $P_{m}=a_{0}^{m}+...+a_{m}^{m}X^{m}$.
Using Equation (3.3) we have
$\Psi_{m}^{\epsilon_{1}}$ $=L(z^{[\epsilon_{1}]_{-}}P_{m}(\chi_{1})z^{[\epsilon_{2}]_{+}})$
$=\sum_{j}a_{j}^{m}L(z^{[\epsilon_{1}]_{-}}\chi_{1}^{j}z^{[\epsilon_{2}]_{+}})$
$=\sum_{j}a_{j}^{m}\Big{(}-jN^{j-1}\alpha-(\mathfrak{p}_{j}-j)N^{j-2}\beta\Big{)}$
$=-\left(\alpha-\frac{\beta}{N}\right)P_{m}^{\prime}(N)-\sum_{j}a_{j}^{m}\mathfrak{p}_{m}N^{j-2}\beta$
$=-\left(\alpha-\frac{\beta}{N}\right)P_{m}^{\prime}(N)-\beta\frac{\mathfrak{p}_{m}}{N^{2}}P_{m}(N).$
Note that we have used the fact that $a_{j}^{m}=0$ whenever $m$ and $j$ are of different parities. Now, using Equation (3.4) we have
$\Psi_{mn}^{\epsilon_{1}\eta_{1}}$ $=L(z^{[\epsilon_{1}]_{-}}P_{m}(\chi_{1})z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}P_{n}(\chi_{1})z^{[\eta_{2}]_{+}})$
$=\sum_{i,j}a_{i}^{m}a_{j}^{n}L(z^{[\epsilon_{1}]_{-}}\chi_{1}^{i}z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}\chi_{1}^{j}z^{[\eta_{2}]_{+}})$
$=\sum_{i,j}a_{i}^{m}a_{j}^{n}\left(-(i+j)N^{i+j-1}\alpha-\Big{(}{(\mathfrak{p}_{i}-\mathfrak{p}_{j})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}^{2}_{ij}-(i+j)}\Big{)}N^{i+j-2}\beta\right)$
$=-\left(\alpha-\frac{\beta}{N}\right)(P_{m}P_{n})^{\prime}(N)-\sum_{i,j}a_{i}^{m}a_{j}^{n}\Big{(}{(\mathfrak{p}_{i}-\mathfrak{p}_{j})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}^{2}_{ij}}\Big{)}N^{i+j-2}\beta$
$=-\left(\alpha-\frac{\beta}{N}\right)(P_{m}P_{n})^{\prime}(N)-\beta\frac{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}}{N^{2}}(P_{m}P_{n})(N).$
Before stating this section’s final theorem, we need one last ingredient. We first need to define the parity entanglement of a tuple $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, to do so let us fix some additional notations.
  • $\ell(\mathbf{n}):=p$ the length of $\mathbf{n}$;
  • $\mathfrak{p}_{\mathbf{n}}:=\sum_{j=1}^{p}\mathfrak{p}_{n_{j}}$ the amount of odd numbers among the $n_{j}$’s;
  • $k_{1}:=\min\{j\geq 1:\mathfrak{p}_{n_{j}}=1\}$ and $k_{i+1}:=\min\{j>k_{i}:\mathfrak{p}_{n_{j}}=1\}$ ($1\leq i<\mathfrak{p}_{\mathbf{n}}$).
We thus define the parity entanglement of $\mathbf{n}$ to be the quantity
$\mathfrak{e}_{\mathbf{n}}:=\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}.$
Lemma 3.5.
Let $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, $\epsilon=\epsilon_{1}\in\{\pm 1\}$ and recursively define $\epsilon_{i+1}:=(-1)^{n_{i}+1}\epsilon_{i}$, then
$\sum_{1\leq i<j\leq p}\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}=2\mathfrak{e}_{\mathbf{n}}.$
Proof.
We may assume that $\mathfrak{p}_{\mathbf{n}}\geq 2$ since both terms are null otherwise. First note that for any $1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}$:
$(-1)^{n_{k_{i}}+...+n_{k_{j-1}}}=(-1)^{j-i}.$
This is because the amount of odd numbers in $\{n_{k_{s}}:i\leq s<j\}$ is exactly $j-i$. This further implies
$\epsilon_{k_{i}}\epsilon_{k_{j}}$ $=(-1)^{n_{k_{i}}+...+n_{k_{j-1}}+k_{j}-k_{i}}\epsilon_{k_{i}}^{2}=(-1)^{n_{k_{i}}+...+n_{k_{j}-1}+k_{j}-k_{i}}=(-1)^{k_{j}+j-(k_{i}+i)}.$
We may now compute
$\sum_{1\leq i<j\leq p}\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}$ $=\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}\delta_{\epsilon_{k_{i}}\epsilon_{k_{j}}}\mathfrak{p}_{n_{k_{i}}n_{k_{j}}}^{2}-\mathfrak{p}_{n_{k_{i}}n_{k_{j}}}$
$=\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}4\delta_{\epsilon_{k_{i}}\epsilon_{k_{j}}}-2$
$=2\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}\epsilon_{k_{i}}\epsilon_{k_{j}}$
$=2\sum_{1\leq j<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}$
$=2\mathfrak{e}_{\mathbf{n}}.$
From the second to third line, we have used the fact that $12\delta_{xy}=xy+$ 1 when $x,y$ are elements in {±$\{\pm 1\}$. Thus, the equality holds. ∎
Theorem 3.6.
Let $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, $\epsilon\in\{\pm 1\}$, then
$L(\chi^{\epsilon}_{\mathbf{n}})=-\left(\alpha-\frac{\beta}{N}\right)P^{\prime}_{\mathbf{n}}(N)-\beta\frac{\mathfrak{p}_{\mathbf{n}}+2\mathfrak{e}_{\mathbf{n}}}{N^{2}}P_{\mathbf{n}}(N),$
where $P_{\mathbf{n}}=P_{n_{1}}...P_{n_{p}}$.
Proof.
Using Lemma A.3, Proposition 3.4 and Lemma 3.5 we have
$L(\chi_{\mathbf{n}}^{\epsilon})$ $=L(z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{[\epsilon_{2}]_{+}}...z^{[\epsilon_{p}]_{-}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}})$
[…]
$\,\,\,\,\,\,\,\,\,\,\,\,\,-(p-1)\sum_{j=1}^{p}\left(-\left(\alpha-\frac{\beta}{N}\right)\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}-\beta\frac{\mathfrak{p}_{n_{j}}}{N^{2}}\right)d_{\mathbf{n}}$
$=-\left(\alpha-\frac{\beta}{N}\right)\sum_{j=1}^{p}\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}d_{\mathbf{n}}-\beta\sum_{1\leq i<j\leq p}\frac{\mathfrak{p}_{n_{i}}+\mathfrak{p}_{n_{j}}-2\mathfrak{p}_{n_{i}}\mathfrak{p}_{n_{j}}+\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}^{2}_{n_{i}n_{j}}}{N^{2}}d_{\mathbf{n}}$
$\,\,\,\,\,\,\,\,\,\,\,\,\,+\beta(p-1)\sum_{j=1}^{p}\frac{\mathfrak{p}_{n_{j}}}{N^{2}}d_{\mathbf{n}}$
$=-\left(\alpha-\frac{\beta}{N}\right)\sum_{j=1}^{p}\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}d_{\mathbf{n}}-\beta\frac{\mathfrak{p}_{\mathbf{n}}}{N^{2}}d_{\mathbf{n}}-\beta\sum_{1\leq i<j\leq p}\frac{\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}}{N^{2}}d_{\mathbf{n}}$
[…]
3.3. Absolute continuity
In Subsection 3.1, we have shown that any Brownian motion on $U_{N}^{+}$ coming from a multiple of the Laplace-Beltrami operator of $U_{N}$ has its constants $\alpha$ and $\beta$ linked by:
$\beta\sim\frac{\alpha}{N},\quad\hbox{as }N\longrightarrow\infty.$
In consequence, we will only consider Brownian motions where $\beta=O(\alpha/N)$. Let $(\psi_{t})_{t\geq 0}$ be a Lévy process associated to such a generating functional, i.e. we have
$\psi_{t}:\chi^{\epsilon}_{\mathbf{n}}\longmapsto d_{\mathbf{n}}\exp\left(-t\left(\left(\alpha-\frac{\beta}{N}\right)\lambda_{\mathbf{n}}+\beta\frac{\mathfrak{p}_{\mathbf{n}}+2\mathfrak{e}_{\mathbf{n}}}{N^{2}}\right)\right),\quad t\geq 0,$
for some $\alpha,\beta\geq 0$ with $\beta=O(\alpha/N)$, where we have written $\lambda_{\mathbf{n}}=P^{\prime}_{\mathbf{n}}(N)/P_{\mathbf{n}}(N)$.
In this paper, to compute limit profiles, we only consider the total variation distance. Just like in the classical setting, the total variation distance for states is much easier to compute when one of the state is absolutely continuous w.r.t the other. A state $\psi$ is said to be absolutely continuous w.r.t the Haar state if there exists $f\in\mathrm{L}^{1}(U_{N}^{+})$ such that $\psi(x)=h(fx)$ for all $x\in\mathcal{O}(U_{N}^{+})$. In this case, the total variation distance is easily expressed (see [21, Lem 2.6])
$d_{\mathrm{TV}}(\psi,h)=\frac{1}{2}\|f-1\|_{1}.$
Quantum groups exhibit a unique characteristic wherein absolute continuity is not easily obtained. Let us first present an elementary fact before stating a proposition on absolute continuity.
Lemma 3.7.
Let $\mathbf{n}\in\bigsqcup_{p\geq 1}\mathbb{N}^{*p}$ be any tuple. We have the following inequality:
$0\leq 2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}\leq\frac{\ell(\mathbf{n})^{2}}{4}.$
Proof.
We will use the notations introduced immediately before Lemma 3.5. Let us set $m_{i}:=k_{i}+i$ ($1\leq i\leq\mathfrak{p}_{\mathbf{n}}$). We further denote by $x$ and $y$ the amount of even and odd numbers among the $m_{i}$’s respectively. We thus have
$2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}$ $=2\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}+x+y$
$=2\left(\frac{x(x-1)+y(y-1)}{2}-xy\right)+x+y$
$=(x-y)^{2},$
this shows the positivity. To show the upper bound, it suffices to notice that if $m_{j}$ and $m_{j+1}$ have the same parity, then we have $k_{j+1}>k_{j}+1$. This further implies that there are at most $\lfloor\ell(\mathbf{n})/2\rfloor$ odd (or even) numbers among the $m_{j}$’s. This allows us to conclude
$2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}=(x-y)^{2}$ $\leq\max(x,y)^{2}\leq\frac{\ell(\mathbf{n})^{2}}{4}.$
We can now characterize absolute continuity.
Proposition 3.8.
Fix $c>0$ and $t=\alpha^{-1}(N\ln(\sqrt{2}N)+cN)$. Provided $N$ is large enough, then $\psi_{t}$ is absolutely continuous w.r.t. the Haar state.
Proof.
We want to see that the series
$f_{t}:=\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}\exp\left(-t\left(\left(\alpha-\frac{\beta}{N}\right)\lambda_{\mathbf{n}}+\beta\frac{\mathfrak{p}_{\mathbf{n}}+2\mathfrak{e}_{\mathbf{n}}}{N^{2}}\right)\right)\chi_{\mathbf{n}}^{\epsilon}$
converges in $\mathrm{L}^{1}(U_{N}^{+})$ for $N$ large enough. We will make use of the $\mathrm{L}^{2}$-norm, as the $\mathrm{L}^{1}$-norm is dominated by it. Take $N$ large enough so that
$\frac{\beta t}{N^{2}}\leq\frac{c}{2}.$
Using the estimates from [17, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have
$\|f_{t}\|_{2}^{2}\leq\sum_{\epsilon,\mathbf{n}}d_{\mathbf{n}}^{2}e^{-2t(\alpha-\beta/N)\lambda_{\mathbf{n}}}\leq 2\sum_{m\geq 1}\sum_{\begin{subarray}{c}\mathbf{n}=(n_{1},...,n_{p}),\\ n_{1}+...+n_{p}=m\end{subarray}}N^{2m}e^{-2t(\alpha-\beta/N)m/N}\leq\sum_{m\geq 1}e^{-cm}<\infty.$
4. Restricting the study to a smaller algebra
Let us recall that the asymptotic study of the Brownian motion on the orthogonal quantum group $O_{N}^{+}$ was done in [23]. However, the proof heavily relies on the commutativity of the central algebra $\mathcal{O}(O_{N}^{+})_{0}$, this is not the case for $\mathcal{O}(U_{N}^{+})_{0}$ (note for instance that $\chi_{1}^{2}\neq\overline{z}\chi_{1}^{2}z$). Our first goal will be to find a smaller commutative subalgebra on which we may restrict our study to drastically ease the computations. We proceed by understanding what would be the smallest algebra $A\subset\mathcal{O}(U_{N}^{+})$ on which all the information of the limit-profile is contained. Before explicitly introducing this algebra, let us present some facts.
Lemma 4.1.
For any tuple $\mathbf{n}$, we have, setting $t_{c}=N\ln(\sqrt{2}N)+cN$ for some $c\in\mathbb{R}$:
$\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})\underset{N\to\infty}{\longrightarrow}e^{-\widetilde{c}|\mathbf{n}|},\quad\hbox{with }\widetilde{c}=c+\ln\sqrt{2}$
Proof.
This follows from the fact that $d_{n_{j}}e^{-t_{c}\lambda_{n_{j}}}\to e^{-\widetilde{c}n_{j}}$ as $N\to\infty$ for any $j\in\{1,...,p\}$ (see the proof of [23, Prop 3.10]). ∎
Now, we understand that the limit-process evaluated on an irreducible character $\chi_{\mathbf{n}}^{\epsilon}$ only depends on $|\mathbf{n}|$, let us introduce compositions. A composition of $m$ is a tuple $\mathbf{n}=(n_{1},...,n_{p})$ that satisfies $|\mathbf{n}|:=n_{1}+...+n_{p}=m$.
Let us denote by $\Pi_{m}$ the set of compositions of an integer $m\in\mathbb{N}^{*}$. Let us point out that $|\Pi_{m}|=2^{m-1}$. We also write
$\Pi_{m}^{1}=\{\mathbf{n}\in\Pi_{m}:n_{1}=1\}\quad\hbox{\&}\quad\Pi_{m}^{>0}=\Pi_{m}\setminus\Pi_{m}^{1}.$
One easily checks that the following maps
$\left\{\begin{matrix}\Pi_{m}&\longrightarrow&\Pi_{m+1}^{1}\\ \mathbf{n}&\longmapsto&(1,\mathbf{n})\end{matrix}\right.,\quad\left\{\begin{matrix}\Pi_{m}&\longrightarrow&\Pi_{m+1}^{>0}\\ \mathbf{n}&\longmapsto&\mathbf{n}^{+}\end{matrix}\right.,$
$\left\{\begin{matrix}\Pi_{m+1}^{1}&\longrightarrow&\Pi_{m}\\ \mathbf{n}&\longmapsto&\mathbf{n}^{-}\end{matrix}\right.\quad\hbox{\&}\quad\left\{\begin{matrix}\Pi_{m+1}^{>0}&\longrightarrow&\Pi_{m}\\ \mathbf{n}&\longmapsto&\mathbf{n}^{-}\end{matrix}\right.,$
are bijections (where we have written $\mathbf{n}^{\pm}=(n_{1}\pm 1,n_{2},...,n_{p})$ and identified $(0,\mathbf{n})\sim\mathbf{n}$).
The last step is identifying a subalgebra that only cares about the composition type. We claim that this algebra is the one generated by
$x:=\frac{\overline{z}\chi_{1}+\chi_{1}z}{\sqrt{2}},$
let us denote it by $\mathcal{O}(U_{N}^{+})_{00}$. First, the restriction of the Haar state coincides with integration w.r.t. the semicircle distribution $\nu_{\mathrm{SC}}$ (see [3, Thm 1]). More precisely, we have $h=\nu_{\mathrm{SC}}\circ\iota$ where $\iota:\mathcal{O}(U_{N}^{+})_{00}\to\mathbb{C}[X]$ is the isomorphism given by $P_{m}(x)=:x_{m}\mapsto P_{m}$. Before stating and proving this section’s final result, let us prove a quick but useful lemma.
Lemma 4.2.
Let $\mathbf{n}$ be any tuple of integer and $\epsilon\in\{\pm 1\}$, we have
$x\cdot\chi_{\mathbf{n}}^{\epsilon}=\frac{\chi_{(1,\mathbf{n})}^{\epsilon}+\chi_{\mathbf{n}^{+}}^{-\epsilon}+\chi_{\mathbf{n}^{-}}^{-\epsilon}}{\sqrt{2}},$
Proof.
Let us write
$\chi_{\mathbf{n}}^{\epsilon}=z^{[\epsilon_{0}]_{-}}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}.$
From there, we simply compute
$\sqrt{2}x\cdot\chi_{\mathbf{n}}^{\epsilon}$ $={(z^{[\epsilon]_{-}}\chi_{1}z^{[\epsilon]_{+}}+z^{[-\epsilon]_{-}}\chi_{1}z^{[-\epsilon]_{+}})z^{[\epsilon_{0}]_{-}}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}}$
$={z^{[\epsilon]_{-}}\chi_{1}z^{\epsilon}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}+z^{[-\epsilon]_{-}}\chi_{1}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}}$
$={\chi_{(1,\mathbf{n})}^{\epsilon}+z^{[-\epsilon]_{-}}(\chi_{n_{1}+1}+\chi_{n_{1}-1})z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}}$
$={\chi_{(1,\mathbf{n})}^{\epsilon}+\chi_{\mathbf{n}^{+}}^{-\epsilon}+\chi_{\mathbf{n}^{-}}^{-\epsilon}}.$
The following proposition shows that for any $m$, the element $x_{m}$ is a linear combination of irreducible characters whose tuples are compositions of $m$. This implies not only that all the information of the limit-process is contained within this algebra but it also gives the existence of a conditional expectation $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{00}$ leaving both the Haar state and the limit-process invariant (seeing as the $x_{m}$’s constitute an orthonormal family for the Haar state).
Proposition 4.3.
The following assertions hold.
  1. (i)
    We have
    $\frac{x_{m}}{2^{m/2}}=\frac{1}{2^{m}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\chi^{\epsilon}_{\mathbf{n}},\quad m\in\mathbb{N}^{*}.$
  2. (ii)
    There exists a conditional expectation $\mathbb{F}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{00}$ leaving the Haar state invariant and such that
    $\mathbb{F}[\chi_{\mathbf{n}}^{\epsilon}]=\frac{\xi_{|\mathbf{n}|}}{2^{|\mathbf{n}|/2}}.$
Proof.
(i) The result is clear for $m=1,2$, assume the result to hold up until some rank $m\geq 2$. To simplify computations, let us write $y_{\mathbf{n}}=\chi^{1}_{\mathbf{n}}+\chi^{-1}_{\mathbf{n}}$. Using Lemma 4.2 we have
$\frac{x_{m+1}}{2^{(m+1)/2}}$ $=\frac{x\cdot x_{m}-x_{m-1}}{2^{(m+1)/2}}$
$=\frac{x}{2^{1/2}\cdot 2^{m}}\sum_{\mathbf{n}\in\Pi_{m}}y_{\mathbf{n}}-\frac{1}{2\cdot 2^{m-1}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}$
$=\frac{1}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m}}y_{(1,\mathbf{n})}+y_{\mathbf{n}^{+}}+y_{\mathbf{n}^{-}}-\frac{1}{2^{m}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}$
$=\frac{1}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m+1}}y_{\mathbf{n}}+\frac{2}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}-\frac{1}{2^{m}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}$
$=\frac{1}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m+1}}y_{\mathbf{n}}.$
This shows the result by induction.
(ii) Let $\mathrm{L}^{2}(U_{N}^{+})_{0}$ and $\mathrm{L}^{2}(U_{N}^{+})_{00}$ be the $\mathrm{L}^{2}$-spaces of $\mathcal{O}(U_{N}^{+})_{0}$ and $\mathcal{O}(U_{N}^{+})_{00}$ respectively and $\mathbb{F}^{\prime}$ be the orthogonal projection $\mathrm{L}^{2}(U_{N}^{+})_{0}\to\mathrm{L}^{2}(U_{N}^{+})_{00}$. Let us show that $\mathbb{F}^{\prime}$ satisfies the relation stated in the proposition, it suffices to show that
$h(\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m})=\frac{\delta_{|\mathbf{n}|,m}}{2^{m/2}},\quad\mathbf{n}\in\bigsqcup_{p\geq 1}\mathbb{N}^{*p},m\in\mathbb{N}^{*}.$
An easy computation enables us to see that the latter relation holds whenever $m\leq 2$. Assume the relation to hold up to some rank $m\geq 2$. Using the traciality of the Haar state and the first item, we have
$h(\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m+1})$ $=h(x\cdot\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m})-h(\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m-1})$
$=\frac{h\big{(}(\chi_{(1,\mathbf{n})}^{\epsilon}+\chi_{\mathbf{n}^{+}}^{-\epsilon}+\chi_{\mathbf{n}^{-}}^{-\epsilon})\cdot x_{m}\big{)}}{\sqrt{2}}-\frac{\delta_{|\mathbf{n}|,m-1}}{2^{(m-1)/2}}$
$=\frac{2\delta_{|\mathbf{n}|+1,m}+\delta_{|\mathbf{n}|-1,m}}{2^{(m+1)/2}}-\frac{\delta_{|\mathbf{n}|,m-1}}{2^{(m-1)/2}}$
$=\frac{\delta_{|\mathbf{n}|,m+1}}{2^{(m+1)/2}}.$
This proves the relation by induction on $m$. It follows from [1, Thm 9.1.2] that $\mathbb{F}=\mathbb{F}^{\prime}\circ\mathbb{E}$, where we have restricted $\mathbb{F}^{\prime}$ to $\mathcal{O}(U_{N}^{+})_{0}$, is the desired conditional expectation. ∎
Finally, let us end this subsection by mentioning that the closure of $\mathcal{O}(U_{N}^{+})_{00}$ in $C(U_{N}^{+})$ is isomorphic to $C([-N,N])$. Indeed, it suffices to see that the spectrum of $x/\sqrt{2}$ is $[-N,N]$. The inclusion $\sigma(x/\sqrt{2})\subset[-N,N]$ follows from the fact that $\|x/\sqrt{2}\|_{C(U_{N}^{+})}\leq N$. On the other hand, there exists a surjective C*-homomorphism $C(U_{N}^{+})\to C(O_{N}^{+}),u_{ij}\mapsto o_{ij}$. In particular, $x/\sqrt{2}$ is sent to $\chi_{1}$ which is well-known to have spectrum $[-N,N]$ (see [9, Lem 4.2]), hence the other inclusion.
5. The limit profile
As stated in the previous section, we would like to restrict the study of the Brownian motion to the subalgebra $\mathcal{O}(U_{N}^{+})_{00}$. Unfortunately, the Brownian motion is not $\mathbb{F}$-invariant, note for instance that
$\psi_{t}(\chi_{2})=(N^{2}-1)e^{-2t(N\alpha-\beta)/(N^{2}-1)}\neq N^{2}e^{-2t(N\alpha+\beta)/N^{2}}=\psi_{t}(\chi_{(1,1)}).$
Let us first determine the limit profile of the process defined by $\widetilde{\psi}_{t}=\psi_{t}\circ\mathbb{F}$.
Theorem 5.1.
We have
$d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\geq d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c<0,$
Proof.
Up to renormalising the process, we may assume that $\alpha=1$. Recall that the process being $\mathbb{F}$-invariant, we may restrict the study to the commutative algebra $\mathcal{O}(U_{N}^{+})_{00}$ for which the $x_{m}$’s form an orthonormal basis.
The proof of the first convergence closely follows the reasoning in the proof [23, Prop 3.10], with a few modifications. Let us outline the elements that differ. Let $c>0$. It follows from Proposition 3.8 that $\widetilde{\psi}_{t_{c}}$ has an $\mathrm{L}^{1}$-density (at least asymptotically defined) given by
$\widetilde{f}_{t_{c}}=\sum_{m\geq 0}{\widetilde{\psi}_{t_{c}}(x_{m})}x_{m}=\sum_{m\geq 0}\frac{1}{2^{m/2}}\left(\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{t_{c}}(\chi^{\epsilon}_{\mathbf{n}})\right)x_{m}.$
Using the isomorphism $\mathcal{O}(U_{N}^{+})_{00}\to\mathbb{C}[X],x_{m}\mapsto P_{m}$ sending the Haar state to the semi-circle distribution, we have
$d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h)=\frac{1}{2}\|\widetilde{f}_{t_{c}}-1\|_{1,h}=\frac{1}{2}\left\|\sum_{m\geq 1}\widetilde{\psi}_{t_{c}}(x_{m})P_{m}\right\|_{1,\nu_{\mathrm{SC}}}.$
In parallel, note that for any tuple $\mathbf{n}=(n_{1},...,n_{p})$, we have $d_{\mathbf{n}}e^{-t_{c}\lambda_{\mathbf{n}}}\to e^{-\widetilde{c}|\mathbf{n}|}$ as $N\to\infty$ where $\widetilde{c}=c+\ln\sqrt{2}$ (this follows from Lemma 4.1 and Proposition 4.3). This further implies that for any given $m\in\mathbb{N}$, we have
$\psi_{t_{c}}(x_{m})$ $=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}|\mathbf{n}|}\exp\left(t_{c}\beta\left(\frac{\lambda_{\mathbf{n}}}{N}-\frac{2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}}{N^{2}}\right)\right)$
$=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}|\mathbf{n}|}\exp\left((N\ln(\sqrt{2}N)+cN)O(N^{-3})\right)=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}m+o(1)}$
$\underset{N\to\infty}{\longrightarrow}2^{m/2}e^{-\widetilde{c}m}=e^{-cm}.$
On the other hand, for any fixed $m\in\mathbb{N}$, we have
$\|\psi_{t_{c}}(x_{m})x_{m}\|_{1}$ $\leq\|\psi_{t_{c}}(x_{m})x_{m}\|_{2}$
$=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{s_{t}}(\chi^{\epsilon}_{\mathbf{n}})$
$\leq\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}(1-\beta/N)\lambda_{\mathbf{n}}}$
$\leq 2^{m/2}\frac{q^{-m}}{(1-q^{2})^{m}}e^{-t_{c}(1-\beta/N)m/N}.$
As mentioned in the proof of Proposition 3.8, the latter ratio can uniformly in $N$ (at least from a certain rank) be bounded by a constant strictly less than one. This implies that we may exchange the sum over $m$ and the limit in $N$, yielding
$\lim_{N\to\infty}d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h)=\frac{1}{2}\left\|\sum_{m\geq 1}e^{-cm}P_{m}\right\|_{1,\nu_{\mathrm{sc}}}=d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}}).$
Now, for the second result, fix $c<0$ and consider the states $\widetilde{\psi}_{t_{c}}$ and $h$ in their measure form as $\mu_{t_{c}}$ and $\nu_{\mathrm{SC}}$. We have already established that $\mu_{t_{c}}$ converges to $\eta_{c}$ in moments as $N\to\infty$. Taking $N$ large enough so that $t_{c}\beta\leq-cN^{2}$, using the estimates from [17, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have for any $m\in\mathbb{N}$
$\mu_{t_{c}}(P_{m})\leq 2^{m/2}N^{m}e^{-t_{c}(1-\beta/N)m/N}\leq e^{-2cm}=\eta_{2c}(P_{m}).$
In particular, this implies that:
$\mu_{t_{c}}(X^{2n})\leq\eta_{2c}(X^{2n}),\quad\text{for all }n\in\mathbb{N}.$
This follows from the fact that $X^{2n}$ can be expressed as a linear combination of $P_{m}$’s with positive coefficients. Since $\eta_{2c}$ is supported on the interval $[-\gamma_{2c},\gamma_{2c}]$, where $\gamma_{2c}=e^{2c}+e^{-2c}$, it follows that $\mu_{t_{c}}$ must also be supported on this interval (as the support of a measure can be estimated by bounding its moments; see, for example, the beginning of [32, Lect 2]). Given that moment convergence with compact support implies weak convergence, we have:
$\mu_{t_{c}}(B)\xrightarrow[N\to\infty]{}\eta_{c}(B),$
for any continuity set $B$ of $\eta_{c}$, i.e., a set such that $\eta_{c}(\partial B)=0$. The final step is to identify such a set that maximizes the total variation distance between $\eta_{c}$ and $\nu_{\mathrm{SC}}$. It suffices to take:
$B_{c}=\{x\in[-2,2]:g_{c}(x)\leq 1\},$
where $g_{c}$ is the density of the absolutely continuous part of $\eta_{c}$ with respect to the semicircular distribution. With this choice of $B_{c}$, we now conclude the argument.
$d_{\mathrm{TV}}(\mu_{t_{c}},\nu_{\mathrm{SC}})\geq\nu_{\mathrm{SC}}(B_{c})-\mu_{t_{c}}(B_{c})\underset{N\to\infty}{\longrightarrow}\nu_{\mathrm{SC}}(B_{c})-\eta_{c}(B_{c})=d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}}).$
We can now conclude for the pure Brownian motion.
Corollary 5.2.
We have
$d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\geq d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c<0,$
where $t_{c}=N\ln(\sqrt{2}N)/\alpha+cN/\alpha$. In particular, the Brownian motion has cutoff at time $t_{N}=N\ln(\sqrt{2}N)/\alpha$.
Proof.
Fix $c>0$, for the first convergence it suffices to see that $d_{\mathrm{TV}}\big{(}\psi_{t_{c}},\widetilde{\psi}_{t_{c}}\big{)}\to 0$ as $N\to\infty$. This is obvious as
$d_{\mathrm{TV}}\big{(}\psi_{t_{c}},\widetilde{\psi}_{t_{c}}\big{)}\leq\sum_{\epsilon,\mathbf{n}}|\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})-\widetilde{\psi}_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})|^{2}\leq\|f_{t_{c}}\|_{2}^{2}+\|\widetilde{f}_{t_{c}}\|_{2}^{2}\leq 2\sum_{m\geq 1}e^{-cm}$
where the last inequality holds for $N$ large enough (similarly to the proof of Proposition 3.8). By dominated convergence, we may exchange summation and limit in $N$ yielding the convergence.
For the second one, simply note that for any $c\in\mathbb{R}$, we have
$d_{\mathrm{TV}}(\psi_{t_{c}},h)=\frac{1}{2}\|\psi_{t_{c}}-h\|_{FS}\geq\frac{1}{2}\|\mathbb{F}\circ\psi_{t_{c}}-\mathbb{F}\circ h\|_{FS}=d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h).$
Let us conclude this section by discussing the result. The limit profile on the right ($c>0$) is straightforward to compute due to the inherent absolute continuity in this region. However, outside of this region, absolute continuity is lost. It is worth recalling that such a loss of absolute continuity was already observed in [23], where all processes studied exhibited an atom as the singular part when absolute continuity was no longer present. P. Biane, in [6, Sec 12.2], introduced an interesting method for recovering the absolutely continuous part of a measure when only its moments are known, provided that one can identify its atom. In our case, the singular part is more complex than a single atom, and it remains unclear whether Biane’s method could be applied here. We conjecture, however, that the singular part lies outside the support of $\nu_{\mathrm{SC}}$, and that there is true $\mathrm{L}^{1}$-convergence. Specifically, we believe that even for $c<0$, the absolutely continuous part of the process converges to that of $\eta_{c}$ in $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$.
6. Non-uniqueness of limit profiles
When examining a limit profile for a family of (quantum) groups, each paired with a central Lévy process $(G_{N},\varphi^{N})_{N\in\mathbb{N}}$, as defined in Definition 2.7, it is pertinent to consider how the profile depends on the structure of $G_{N}$ and the dynamics of $\varphi^{N}$. Although the impact of the group’s representation theory on the profile is obvious (see the profiles studied in [23, 30, 33] and this paper), the degree to which the process itself influences the limit profile is not clear. This raises a question: is the temporal aspect the only significant parameter through which the process affects the cutoff behavior?
This section presents an example to demonstrate that different central Lévy processes on the same quantum group can lead to distinct limit profiles. Consider the orthogonal quantum group $O_{N}^{+}$ and its associated Brownian motion, which typically has a limit profile function $c\mapsto d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}})$. We define a new Lévy process on $O_{N}^{+}$, designed to exhibit a different limit profile.
The central generating functional for this new process is defined by:
$L:\chi_{n}\mapsto-P^{\prime}_{n}(N)-\frac{P_{n}(N)-P_{n}(-N)}{2}=\left\{\begin{array}[]{ll}-P_{n}^{\prime}(N)&\text{if }n\text{ is even},\\ -P_{n}^{\prime}(N)-P_{n}(N)&\text{otherwise}.\end{array}\right.$
This is the functional with parameters $b=1$ and $\nu=N\delta_{-N}$, when considering the decomposition in Equation (3.1). The resulting Lévy process $\varphi_{t}$ is then given by:
$\varphi_{t}:\chi_{n}\longmapsto d_{n}e^{-t\lambda_{n}}c_{t,n}\quad\text{where }c_{t,n}=\left\{\begin{array}[]{ll}1&\text{if }n\text{ is even},\\ e^{-t}&\text{otherwise}.\end{array}\right.$
Theorem 6.1.
Fix $c\in\mathbb{R}$ and set $t_{N}:=N\ln(N)+cN$, we have
$d_{\mathrm{TV}}(\varphi_{t_{N}},h)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}(\hat{\eta}_{c},\nu_{\mathrm{SC}}),$
where $\hat{\eta}_{c}$ is the unique distribution satisfying $\hat{\eta}_{c}(P_{n})=e^{-nc}$ if $n$ is even and 00 otherwise. The distribution $\hat{\eta}_{c}$ can be described by:
$\hat{\eta}_{c}=\mathbf{1}_{c\leq 0}(1-e^{2c})\frac{\delta_{\gamma_{c}}+\delta_{-\gamma_{c}}}{2}+\frac{e^{c}\gamma_{c}}{\gamma_{c}^{2}-x^{2}}\mathrm{d}\nu_{\mathrm{SC}}(x),\quad\text{with }\gamma_{c}=e^{c}+e^{-c}.$
Proof.
Fix $c\in\mathbb{R}$. Let us first prove that the distribution $\hat{\eta}_{c}$ as described in the theorem has the desired moments. If $c>0$, then clearly the series $\sum_{n\geq 0}e^{-2cn}P_{2n}$ belongs to $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$. Applying twice the recursive formula satisfied by the Chebyshev polynomials confirms that the series is the density of $\hat{\eta}_{c}$ w.r.t $\nu_{\mathrm{SC}}$. If $c=0$, then we conclude by dominated convergence
$\hat{\eta}_{0}(P_{n})=\int_{-2}^{2}\frac{2P_{n}(x)}{2-x^{2}}\mathrm{d}\nu_{\mathrm{SC}}(x)=\lim_{c\to 0^{+}}\hat{\eta}_{c}(P_{n})=\left\{\begin{array}[]{cl}1&\hbox{if }n\hbox{ is even}\\ 0&\hbox{else},\end{array}\right.\quad n\in\mathbb{N}.$
Finally, if $c<0$, using the positive case, we have whenever $n$ is even
$\hat{\eta}_{c}(P_{n})$ $=(1-e^{2c})\frac{P_{n}(\gamma_{c})+P_{n}(-\gamma_{c})}{2}+e^{2c}\hat{\eta}_{-c}(P_{n})$
$=(1-e^{2c})\frac{e^{-(n+1)c}-e^{(n+1)c}}{e^{-c}-e^{c}}+e^{(n+2)c}=e^{-nc}.$
Clearly, the computation leads to zero when $n$ is odd.
Now to prove the limit profile recall that we can see $\varphi_{t_{N}}$ and $h$ as their measure counterparts $\mu_{t_{N}}$ and $\nu_{\mathrm{SC}}$ via the isomorphism $\chi_{n}\mapsto P_{n}$. It remains to see that the absolutely continuous part of $\mu_{t_{N}}$ w.r.t $\nu_{\mathrm{SC}}$ tends to that of $\hat{\eta}_{c}$ as $N\to\infty$ in total variation distance which can be done in a similar fashion as in the proof of [23, Thm 3.9]. ∎
Observe that the functions $f_{1}:c\mapsto d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}})$ and $f_{2}:c\mapsto d_{\mathrm{TV}}(\hat{\eta}_{c},\nu_{\mathrm{SC}})$ can be explicitly computed through integration, resulting in the following expressions:
$f_{1}(c)=\begin{cases}\frac{2(1-e^{2c})}{\pi}\arcsin\left(\frac{e^{-c}}{2}\right)+\frac{e^{-2c}+2}{4\pi}\sqrt{4e^{2c}-1},&\text{if }c\geq-\ln 2,\\ 1-e^{2c},&\text{if }c<-\ln 2,\end{cases}$
and
$f_{2}(c)=\begin{cases}\frac{\sqrt{1+e^{-2c}}\sqrt{3-e^{-2c}}}{2\pi}+\frac{e^{2c}-1}{\pi}\left(\arctan\left(\frac{\sqrt{1+e^{-2c}}(e^{c}-e^{-c})}{(e^{c}+e^{-c})\sqrt{3-e^{-2c}}}\right)-\arcsin\left(\frac{\sqrt{1+e^{-2c}}}{2}\right)\right),&\text{if }c\geq-\frac{\ln 3}{2},\\ 1-e^{2c},&\text{if }c<-\frac{\ln 3}{2}.\end{cases}$
Figure 1. Graph comparing the functions $f_{1}$ (red) and $f_{2}$ (blue).
They agree on infinitely many points but are indeed distinct. This implies that they are not equivalent modulo affine transformation. It follows from Remark 2.8 that the profile we exhibited is different.
Interestingly, although we might have anticipated that the new process would result in a ’faster shuffle’, in the sense that $f_{2}\leq f_{1}$, the two profiles initially overlap for a significant range before the impact of the ’faster shuffle’ becomes apparent.
Appendix
In this appendix, we give a description of Gaussian centralized generating functionals on $O_{N}^{+}$ and $U_{N}^{+}$. We will therefore focus our study on the latter quantum groups, we refer the reader to [19, Sec 1.5] for definitions in greater generality.
We would like to acknowledge that this result was previously achieved by different means by Franz, Freslon, and Skalski in an unpublished work providing a comprehensive description of Gaussian generating functionals on certain quantum groups.
Definition A.1.
Let $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$ and $D$ a pre-Hilbert space. We call Schürmann triple on $\mathbb{G}$ over $D$ a family of linear maps $(\rho,\eta,L)$ where
[…]
  • $\eta:\mathcal{O}(\mathbb{G})\to D$ satisfies
    $\eta(ab)=\rho(a)\eta(b)+\eta(b)\varepsilon(b),\quad a,b\in\mathcal{O}(\mathbb{G});$
  • $L:\mathcal{O}(\mathbb{G})\to\mathbb{C}$ is hermitian and satisfies
    $L(ab)=\varepsilon(a)L(b)+L(a)\varepsilon(b)+\langle\eta(a^{*}),\eta(b)\rangle,\quad a,b\in\mathcal{O}(\mathbb{G}).$
The definition implies that $L$ is a generating functional. Two Schürmann triples $(\rho_{i},\eta_{i},L)$ ($i=1,2$) on $\mathbb{G}$ and over pre-Hilbert spaces $D_{i}$ are said to be equivalent, if there exists a surjective isometry $U:D_{1}\to D_{2}$ s.t.
$\eta_{2}(a)=U\eta_{1}(a)\rho_{2}(a)U=U\rho_{1}(a),\quad a\in\mathcal{O}(\mathbb{G}).$
There is a one-to-one correspondence between Schürmann triples with surjective cocycle (up to equivalence) and generating functionals.
Proposition A.2.
Let $L$ be a generating functional on $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$. Then the following are equivalent
  1. (i)
    $L(abc)=L(ab)\varepsilon(c)+L(ac)\varepsilon(b)+L(bc)\varepsilon(a)-\varepsilon(ab)L(c)-\varepsilon(ac)L(b)-\varepsilon(bc)L(a)$ for all $a,b,c\in\mathcal{O}(\mathbb{G})$;
  2. (ii)
    $\eta(ab)=\varepsilon(a)\eta(b)+\eta(a)\varepsilon(b)$ for all $a,b\in\mathcal{O}(\mathbb{G})$.
If a Schürmann triple satisfies one of these conditions, then we call it Gaussian. The Gaussian property of a generating functional gives a way to compute its values on any product of elements.
Lemma A.3.
[16, Prop 2.7] Let $L$ be a Gaussian generating functional on $\mathbb{G}=O_{N}^{+}$ or $U_{N}^{+}$ Then, for any $a_{1},...,a_{n}\in\mathcal{O}(\mathbb{G})$ we have
$L(a_{1}...a_{n})=\sum_{1\leq i<j\leq n}L(a_{i}a_{j})\varepsilon(a_{0}...\check{a}_{i}...\check{a}_{j}...a_{n})-(n-2)\sum_{j=1}^{n}L(a_{j})\varepsilon(a_{1}...\check{a}_{j}...a_{n}).$
Proposition A.4.
We have the following decompositions.
  • Any centralized Gaussian generating functional on $O_{N}^{+}$ is of the form
    $\chi_{n}\longmapsto-bP_{n}^{\prime}(N)$
    for some $b\geq 0$.
  • Any centralized Gaussian generating functional on $U_{N}^{+}$ is of the form
    $\chi_{w}\longmapsto-{\ell(w)}N^{\ell(w)-1}\alpha-\Big{(}{(p(w)-q(w))^{2}-\ell(w)}\Big{)}N^{\ell(w)-2}\beta,\quad w\hbox{ word on }\ \{\lozenge,\blacklozenge\},$
    for some $\alpha\in\mathbb{C}$ and $\beta\geq 0$ satisfying $\mathrm{Re}(\alpha)\geq\beta/N$.
Proof.
Let us first note that any Gaussian cocycle $\eta$ defined on a compact quantum group satisfies the relation $\eta\circ S+\eta=0$ (see the beginning of the proof of [16, Thm 3.11]).
Let $(\rho,\eta,L)$ be a Gaussian Schürmann triple on $O_{N}^{+}$. The relation $\eta\circ S=-\eta$ implies that $\eta\equiv 0$ on the central algebra and therefore, we have
$L(ab)=\varepsilon(a)L(b)+L(a)\varepsilon(b),\quad a,b\in\mathcal{O}(O_{N}^{+})_{0}.$
It follows from a quick induction that $L$ is of the required form for $b:=-L(\chi_{1})$.
Let $(\rho,\eta,L)$ be a Gaussian Schürmann triple on $U_{N}^{+}$. First note that the relation only needs to be proven for characters $\chi_{w}$ where $w$ is a word of length $\leq 2$ as the general formula would then follow from Lemma A.3. The relation $\eta\circ S=-\eta$ implies that $\eta+\overline{\eta}\equiv 0$ on the central algebra and therefore, we have
$L(ab)=\varepsilon(a)L(b)+L(a)\varepsilon(b)-\langle\eta(a),\eta(b)\rangle,\quad a,b\in\mathcal{O}(U_{N}^{+})_{0}.$
It follows from a quick computation that $L$ is of the required form for $\alpha:=-L(\chi_{\circ})$ and $\beta:=\|\eta(\chi_{\circ})\|^{2}/2$. ∎
We conclude by explaining why we restrict our study to real-valued (centralized) Gaussian generating functionals on $U_{N}^{+}$. If we were to consider the imaginary part of $\alpha$, convergence may not be guaranteed unless a bound exists that links the real and imaginary parts, similarly to the relation between $\alpha$ and $\beta$. However, we emphasize that, by introducing a drift to the generating functional, the imaginary part can always be disregarded.
Remark A.5.
A drift on $U_{N}^{+}$ is defined as a generating functional whose associated cocycle is zero. In other words, a drift is an $\varepsilon$-derivation, i.e., a functional $D:\mathcal{O}(U_{N}^{+})\to\mathbb{C}$ satisfying
$D(ab)=\varepsilon(a)D(b)+D(a)\varepsilon(b),\quad a,b\in\mathcal{O}(U_{N}^{+})_{0}.$
Note that a drift is completely determined by its values on the generators. More precisely, there is a Lie algebra isomorphism between the algebra of drifts and the Lie algebra $\mathfrak{u}_{N}$ of the classical group $U_{N}$. For each skew-Hermitian matrix $H\in\mathfrak{u}_{N}$, we associate a drift $D_{H}$ by setting $D_{H}(u_{ij})=H_{ij}$. The Lie bracket is given by
$[D_{H},D_{K}]=D_{H}\star D_{K}-D_{K}\star D_{H}=D_{[H,K]}.$
This shows that drifts correspond to the classical part of Gaussian generating functionals. By adding such a drift, we can always assume a centralized Gaussian generating functional to be real-valued.